Two pumps connected in parallel fail independently of one another on any given day. The probability that only the older pump will fail is 0.25, and the probability that only the newer pump will fail is 0.15. What is the probability that the pumping system will fail on any given day (which happens if both pumps fail)? (Round your answers to four decimal places.)

I had put 1/4 and 1/2 however these were proven to be incorrect?
Answer #1

P(older pump fail)=0.25

P(new pump fail) = 0.15

P(pumping system will fail )=P(both pumps fail) = P(older pump fail)*P(new pump fail) = 0.25*0.15 = 0.0375

---------------------------

please revert if have doubts